A sphere is charged to 20 µC. Find the potential at the following points. A, B and C (A point is at the center of the sphere, B is at the surface and C is 4cm from the surface.) AB r1=2cm, BC r2=4cm


1
Expert's answer
2020-08-10T19:53:09-0400

"V_A=\frac{(9\cdot10^{9})(20\cdot10^{-6})}{0.02}=9\cdot10^{6}V\\V_B=\frac{(9\cdot10^{9})(20\cdot10^{-6})}{0.02}=9\cdot10^{6}V\\V_C=\frac{(9\cdot10^{9})(20\cdot10^{-6})}{0.02+0.04}=3\cdot10^{6}V"